what volume of .100 M NaOH is needed to make 100.0mL of a buffer solution with a pH of 6.00 if one starts with 50.0mL of 0.100M potassium hydrogen phthalate? The Ka2 for potassium hydrogen phthalate is 3.1x10^-6
Answer #1


aud pH = pka+log ( salt) 6.00 = 5.51 +log ( salt) log (salt) = 6.00 -515) > 0.49 (cale = 3:09 moles of acid = Mxv =0.1000x0.0